Solve (2x - 3)2 = 49 by using the Square Root Method
Answer by Alan3354(69443)

Solve (2x - 3)2 = 49 by using the Square Root Method
-----------
Use ^ (Shift 6) for exponents.
(2x - 3)^2 = 49
---
2x - 3 = ±7
---
2x - 3 = 7
x = 5
====
2x - 3 = -7
x = -2